write an equation of the line that is parallel to the given line and passes through the given point.y=-3x + 2,(2,3)

Answer by robertb(5830)

The line has to have the equation
y = -3x + c, where c is an unknown constant. To find its value, plug in the coordinates of the given point: 3 = -3*2 + c, or 3 = -6 + c, or c = 9.
Therefore the equation is y = -3x + 9.